OpenStudy (samsan9):

Solve the system of equations using either the substitution method or the multiplication/addition method answer: 3x+2y=14 2x-4y=46x+4y=28 2x-4y=4 8x = 32 x = 4 2y = 14 - 12 = 2 y = 1 (x , y) = (4 , 1) question: Check your solution by writing the system as a matrix equation and using the inverse matrix. (how do I do that?)

OpenStudy (samsan9):

my question is how do you turn them into a matrixand then doing using the inverse matrix.

hero (hero):

\(Ax = B\) \(x =A^{-1}B\) \[A = \left[ \begin{matrix} 3 & 2 \\ 2 & 4 \end{matrix} \right]\]\[B = \left[ \begin{matrix} 14 \\ 4 \end{matrix} \right]\] \[\left[ \begin{matrix} 3 & 2 \\ 2 & 4 \end{matrix} \right] \left[ \begin{matrix} x \\ y \end{matrix} \right] = \left[ \begin{matrix} 14 \\ 4 \end{matrix} \right]\] Find the inverse of matrix A, then multiply it by matrix B

OpenStudy (samsan9):

so do you just do it in reverse sorry im lost :(

OpenStudy (samsan9):

how do you find the inverse?

OpenStudy (triciaal):

what are your original equations and did you already solve using the substitution method or the multiplication/addition method ? sorry but it is not very easy to understand your post; everything is all pushed together. Do you only need help with checking using the matrix method?

OpenStudy (samsan9):

my question is this Check your solution by writing the system as a matrix equation and using the inverse matrix

OpenStudy (samsan9):

i already solve my first question and i got (1,4) and the numbers were 3x+2y=14 and 2x-4y=4

OpenStudy (triciaal):

ok thanks

OpenStudy (samsan9):

it is not easy huh?

OpenStudy (triciaal):

I would have to review matrices

OpenStudy (samsan9):

http://www.youtube.com/watch?v=iH3CUVI9S4k is this helpful ?

hero (hero):

\(A^{-1} = \dfrac{1}{|A|} \left[ \begin{matrix} d & -b \\ -c & a \end{matrix} \right]\)

OpenStudy (triciaal):

same as @Hero and ad-bc is not = 0

OpenStudy (samsan9):

is it